[Hallc_running] [New Logentry] RC Daily Update

bduran at jlab.org bduran at jlab.org
Fri Aug 28 13:35:02 EDT 2020

Logentry Text:
Yesterday we had an eventful ~ 2 hour access (11am - 12:40pm) and we completed the following tasks:

 1) HV supply for the transport and compton ion chamber is separated to troubleshoot the ion chamber issue (https://logbooks.jlab.org/entry/3835254), Jerry Kowal and his team was in charge of this (https://logbooks.jlab.org/entry/3835705).
 2) Joe Beaufait resetted the rotation drive controllers. 
 3) Stan, Jerry and Larry connected the spare air compressor that Walter rented. (https://logbooks.jlab.org/entry/3833255, https://logbooks.jlab.org/entry/3835706)
 4) Junhao changed the EPR RF amplifier gain setting from 70% to 40%  to fix  the noisy NMR (https://logbooks.jlab.org/entry/3835639) (https://logbooks.jlab.org/entry/3835639)
 5) Brad checked if the blowers are on to troubleshoot (https://logbooks.jlab.org/entry/3835710 https://logbooks.jlab.org/entry/3835710, https://logbooks.jlab.org/entry/3835740, https://logbooks.jlab.org/entry/3835740) 

At 12:40 pm, we were back in beam permit. After checking the beam positions and possible scraping on reference cell (He3)  (https://logbooks.jlab.org/entry/3835696, https://logbooks.jlab.org/entry/3835702) we resumed data taking with this target. An NMR @2:08pm gave 35.8% https://logbooks.jlab.org/entry/3835727 and pNMR: 35.27%  https://logbooks.jlab.org/entry/3835735. After consulting with JP and Brad, we decided to start data taking with PolHe3 with transverse spin. After verifying the beam positions and no scraping (https://logbooks.jlab.org/entry/3835752, https://logbooks.jlab.org/entry/3835792), we took data with PolHe3 transverse until the accelerator went down at 9:35pm due to weather-caused RF trips. At 7:38pm another NMR/pNMR measurement was performed, 35.52/35.8% respectively (https://logbooks.jlab.org/entry/3835936, https://logbooks.jlab.org/entry/3835936). At 8:44pm, Bill started transverse QWP optimization (https://logbooks.jlab.org/entry/3835943). The beam qual!
 ity during the data taking wasn't great due to mostly BLM trips at injector/HallD line and occasional Hall A/C ion chamber trips. During the down time, as a follow up to the transport ion chamber work done in the hall yesterday, Hall C ion chamber HV was power cycled (https://logbooks.jlab.org/entry/3835989). We will see if this will help with the ion chamber trip problem.

MCC was ready to send beam  at 1:07am. After a long downtime, we asked for harp scans to check the beam profile ( https://logbooks.jlab.org/entry/3836043, https://logbooks.jlab.org/entry/3836053) and  following that, checked the beam positions for PolHe3 at no target (https://logbooks.jlab.org/entry/3836067). Before we started running with PolHe3 another NMR/pNMR measurement was done @1:45am: 37.4%/37.2%   (https://logbooks.jlab.org/entry/3836071 https://logbooks.jlab.org/entry/3836066). When scraping check was done at 1:47am, we resumed data taking with PolHe3 transverse. We took data until 05:10am when SL RF went down because of iocch2a being unresolved (https://logbooks.jlab.org/entry/3836159), this was followed by SL 2K cold box trip as it's unable to handle heat load changes with only half the SL up. 

The accelerator has been down since 05:10am. At today's MCC morning meeting, the expected down time was expected as at least 5-6 hours. We had an update at 10:57am: they were able to reboot and restore IOCCH2A and regain control over the SW linac signals. There is an LHe overflow from SL return and they're setting up the system for the overflow recovery which will likely won't happen before tomorrow morning (https://logbooks.jlab.org/entry/3836302). Fairly soon we had another update 11:07am indicated the situation doesn't look as bad as they initially thought. However, based on my conversation with crew chief this still means a couple of hours down time even if the pumpdown attemp is successful (https://logbooks.jlab.org/entry/3836308). At 12:18pm, the pumpdown attempts have started https://logbooks.jlab.org/entry/3836334.

Melanie's AFP loss test on Briana at 270 deg @10:13am : https://logbooks.jlab.org/entry/3836289

At 10:51am today, Junhao took an escorted access to test the impact of the EPR RF amplifier on NMR signal (https://logbooks.jlab.org/entry/3836301, https://logbooks.jlab.org/entry/3836329).  Since the expected downtime is a couple of more hours and we'd like to take advantage of it for Simona's SoLID Cherenkov work and Junhao's possible EPR RF test we asked for a full survey which was finished at 12:18pm. Simona is in the hall for her SoLID Cherenkov work.

Target temperature test @1:31pm : https://logbooks.jlab.org/entry/3836349

We will have an RC Meeting at 4:15 pm

Bluejeans remote connection information:
To join the Meeting:
To join via Room System:
Video Conferencing System: bjn.vc -or-
Meeting ID : 861439920
To join via phone :
1) Dial: +1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3836354
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20200828/719fe0f9/attachment.html>

More information about the Hallc_running mailing list